OpenStudy (anonymous):

Factoring By Grouping Factor each completely. If non-factorable, write "Non-factorable"

OpenStudy (anonymous):

r^3 - 8r^2 + 5r - 40

Directrix (directrix):

r^3 - 8r^2 + 5r - 40 = r^2 ( r - 8) + 5 (r - 8) = Note: (r - 8) is a common factor ( r - 8) ( ? + ) @x_n3rd_x